# Co-Reinforcement Learning for Unified Multimodal Understanding and Generation

### v2

@@ 1 Introduction
As large foundation models (LFMs) continue to advance in general capabilities and breadth of knowledge, post-training [27, 67, 101] has emerged as a critical paradigm for further refining pretrained LFMs toward specialized applications, facilitating task adaptation and human-aligned behaviors. Recently, reinforcement learning (RL)-based approaches [47, 48, 66, 88, 64, 55, 60] have exhibited considerable promise due to their data efficiency and strong alignment abilities. A notable exemplar is DeepSeek-R1 [18], which demonstrates that RL with verifiable rewards and the group relative policy optimization (GRPO) algorithm constitutes a practical and stable strategy that sidesteps explicit preference modeling [70] and reward model learning [76]. This promising paradigm indicates the significant potential of LFMs to acquire advanced capabilities and generalize effectively without dependence on large-scale, high-quality supervised data.
@@
In the multimodal AI research community, the prevailing implementation [37, 22, 8, 49, 38, 94, 93, 64] of the GRPO algorithm centers on crafting diverse rule-based reward mechanisms to incentivize long-chain reasoning capabilities of multimodal large language models (MLLMs). These initiatives primarily target multimodal understanding, with particular focus on visual and mathematical reasoning tasks. Conversely, its application to visual generation remains surprisingly limited, with only pioneering explorations [72, 23] suggesting its feasibility. More importantly, extending GRPO to unified MLLMs (ULMs) [7, 83, 81, 36] capable of concurrently performing visual understanding and generation tasks remains considerably under-explored. Intuitively, ULMs could significantly benefit from GRPO owing to their inherent advantages of cross-task synergy and LLM sharing, which enables ULMs to share reward signals across various tasks and effectively mitigate reward imbalance, particularly when GRPO operates by jointly ranking outputs within task-agnostic groups.
@@
This paper aims to enhance the understanding and generation capabilities of ULMs without dependence on supervised data. We begin with pilot experiments to explore efficient reinforcement learning paradigms. Specifically, we systematically examine four rule-based training strategies: (i) separate RL for individual tasks, (ii) separate RL with weight merging, (iii) cycle RL alternating between tasks, and (iv) unified RL with joint optimization. Our explorations reveal two critical findings. First, direct task-specific RL fails to achieve the anticipated improvements, particularly in visual generation, and even impairs other abilities. Second, compared with alternative strategies, unified RL showcases comprehensive advantages across tasks. These results demonstrate the synergistic co-evolution of dual capabilities under a shared policy optimization paradigm.
@@
In light of our preliminary findings, we propose CoRL, a co-reinforcement learning framework designed to synergistically improve the understanding and generation capabilities of ULMs. CoRL follows a two-stage RL procedure: a unified RL stage for joint optimization of dual capabilities and a refined RL stage for task-specific enhancement. In the first stage, the policy ULM is optimized through a unified GRPO algorithm with diverse rewards on a carefully curated dataset spanning both understanding and generation tasks. To effectively guide policy optimization in visual generation, we introduce a bidirectional cycle consistency reward and a text-image matching reward, which together promote semantic consistency and faithfulness of synthesized images to their corresponding prompts. The designed rewards complement typical multimodal understanding rewards (i.e., accuracy and format) within a unified group, enabling cross-task joint optimization. In the subsequent stage, we independently reinforce the policy’s understanding and generation capabilities using respective rewards and tailored datasets for task-specific refinement.
@@
Applying the two-stage CoRL training to the baseline ULM Janus-Pro [7] yields ULM-R1, a unified model with reinforced capabilities in both understanding and generation. To comprehensively assess its performance, we conduct extensive comparisons against state-of-the-art unified MLLMs and dedicated models across three visual generation and nine multimodal understanding benchmarks. Notably, ULM-R1 achieves substantial gains over its baseline on complex mathematical and logical reasoning tasks, such as WeMath (+15.2) and LogicVista (+10.6). These results underscore the effectiveness of CoRL, providing compelling empirical evidence for the efficacy of RL in simultaneously advancing visual understanding and generation tasks.
@@ 2 Related Work
Unified Multimodal Understanding and Generation.  Recent advancements [80, 7, 36, 81, 79, 65, 83, 100, 84, 75, 62, 69] have witnessed increasing attention to jointly model multimodal understanding and visual generation within a unified model. Pioneering attempts [16, 11] predominantly rely on continuous diffusion models, integrating external diffusion decoders for image synthesis. Inspired by autoregressive next-token prediction, a growing line of research [65, 80, 75, 78, 7, 53, 36, 81, 79, 28, 43] encode visual inputs into discrete tokens and generate images in a fully autoregressive (F-AR) manner. Specifically, this approach employs a vector quantized (VQ) tokenizer [13, 86] to convert images into discrete tokens, analogous to text tokenization. To mitigate information loss in VQ discretization, another stream of work [83, 69, 21, 26, 15, 44, 85] explores autoregressive and diffusion (AR-Diff) hybrid modeling approaches. Architecturally, these models typically comprise a vision autoencoder, a text tokenizer, and an LLM. This work builds upon the F-AR model to develop an effective reinforcement learning framework.
@@
RL-based Post-Training for MLLMs.  Post-training [101] aims to further enhance the performance of pretrained models for customized applications and user needs. Recently, RL [77, 71] has emerged as a powerful post-training technique, enabling models to learn from feedback and align with human values. RL in MLLMs can be broadly categorized into two paradigms: (1) RL from human/AI feedback (RLHF) [75, 85, 76, 63, 88, 89, 102, 31, 99, 50, 70, 57, 74, 92] and (2) RL with verifiable reward mechanisms [72, 93, 64, 38, 37, 32]. RLHF involves learning reward models from preference data before RL optimization, whereas the latter directly optimizes models using task-specific reward functions, bypassing explicit preference modeling. For example, DPO [55] is a notable implementation of RLHF and has been adopted by Emu3 [75] and HermesFlow [85] to narrow the performance gap between understanding and generation. In contrast, GRPO [60] exemplifies the second paradigm, simplifying reward formulation via group-wise relative advantage estimation. Our work also falls into this paradigm but diverges from prior work such as SimpleAR [72], which utilizes GRPO with external CLIP reward for autoregressive visual generation, and R1-like MLLMs [93, 22, 64, 38] that focus on incentivizing reasoning capabilities. First, our work demonstrates the significant potential of RL in co-optimizing understanding and generation, thereby broadening its applicability beyond reasoning. Moreover, we identify semantic consistency rewards and a co-evolutionary reinforcement strategy as crucial components in enhancing ULMs.
@@ 3.1 Preliminary
Group relative policy optimization (GRPO) [60] is a value-free policy optimization algorithm with improved training stability and sample efficiency. Building upon PPO [58], GRPO introduces a group-wise relative advantage approach to bound policy updates while maintaining optimization flexibility. Let $\pi_{\bm{\theta}}$ denote a policy parameterized by $\bm{\theta}$. Formally, given an input content $c$, the algorithm first samples a group of $G$ outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the current policy $\pi_{\bm{\theta}_{\mathrm{old}}}$. Each output is then evaluated using predefined, verifiable reward functions, yielding the reward set $\{r_{1},r_{2},\dots,r_{G}\}$. These rewards are subsequently normalized to compute group-relative advantages as follows:
@@ 3.2 Pilot Exploration
Figure 1: Results of different RL paradigms. Janus-Pro-1B [7] serves as the baseline.
Given the exceptional performance and data efficiency of DeepSeek-R1-Zero [18], we explore the potential of ULMs to enhance understanding and generation capabilities without dependence on task-specific supervised fine-tuning. To accomplish this, we curate a dataset [^1] https://huggingface.co/datasets/mm-vl/x2x_rft_16k comprising 16K samples sourced from the COCO 2017 train split [35]. Each sample includes a real image, an associated caption as a textual prompt for visual generation, and a corresponding QA pair for the multimodal understanding task. We adopt CLIP Score [54] as the verifiable reward for image generation, and a combination of formatting correctness and answer accuracy as the reward for text generation. We investigate four distinct RL paradigms: (i) separate RL, where understanding and generation tasks are independently optimized under their respective reward mechanisms; (ii) separate RL followed by weight merging, where each task is separately optimized, and the resulting weights are subsequently merged using a Gaussian distribution-based merging strategy [61] to incorporate both abilities; (iii) cycle RL, which applies a scheduled alternation between the two tasks throughout the training process; and (iv) unified RL, in which both tasks are jointly optimized within a unified paradigm to promote co-evolution.
As presented in Figure 1, we observe that (1) direct task-specific RL fails to achieve the expected improvements for ULMs, particularly in the visual generation task, and may even impair performance on the other task; and (2) unified RL demonstrates substantial advantages over alternative paradigms. These findings indicate that the dual capabilities co-evolving within a shared training framework contribute to more effective cross-task synergy and knowledge transfer.
@@ 3.3.1 Verifiable Reward for Multimodal modeling
Bidirectional Cycle Consistency Reward in Text-to-Image Generation.  To encourage ULMs to generate images that faithfully reflect the concepts and entities described in the input prompt, we introduce a bidirectional cycle consistency reward $\mathcal{R}_{\text{cycle}}$, which measures the consistency between predictions and ground truth in both visual and textual spaces. For visual consistency, we adopt LPIPS [97] to assess the patch-level perceptual similarity between the real image $\mathcal{I}_{\text{real}}$ and the synthesized image $\mathcal{I}_{\text{gen}}$. Textual consistency is implemented in a re-captioning manner. Specifically, we first employ BLIP [29] to generate a caption $\mathcal{C}_{\text{re-cap}}$ for each synthesized image, and then compute the SPICE [1] score between $\mathcal{C}_{\text{re-cap}}$ and its original prompt $\mathcal{P}_{\text{org}}$ to measure semantic fidelity. The combined bidirectional cycle reward is defined as:
@@
This bidirectional reward forms a closed feedback loop that promotes mutual consistency between text and image, effectively penalizing hallucinated content and reinforcing prompt-adherent visual generation by simultaneously optimizing for both visual and textual consistency.
Text-Image Matching Reward.  While CLIP Score [54] provides a holistic measure of text-image alignment, it underperforms in Sec. 3.2 due to its limited ability to assess fine-grained semantics. To address this limitation, we instead propose a text-image matching reward $\mathcal{R}_{\text{TIM}}$, which leverages the ULM itself to assess cross-modal alignment at the token level. Given a textual representation $\bm{T}=\{\bm{t}_{1},\bm{t}_{2},\ldots,\bm{t}_{L_{t}}\}\in\mathbb{R}^{L_{t}\times d}$ of the prompt and the corresponding visual representation $\bm{I}=\{\bm{i}_{1},\bm{i}_{2},\ldots,\bm{i}_{L_{i}}\}\in\mathbb{R}^{L_{i}\times d}$ of a generated image, the reward is computed as:
@@
where $L_{t}$ and $L_{i}$ are the lengths of textual and visual token sequences, and $d$ is the embedding dimension. This reward captures the fine-grained correspondence between textual concepts and visual elements through maximum cosine similarity, ensuring that each visual token aligns with its most relevant textual counterpart and vice versa.
Accuracy Reward in Multimodal Question Answering.  Accuracy rewards leverage task-specific metrics to directly evaluate the correctness of ULM predictions. We consider two accuracy rewards tailored to different question types: $\mathcal{R}_{\text{MCQ-Acc}}$ for multi-choice questions and $\mathcal{R}_{\text{OE-Acc}}$ for open-ended questions. These rewards follow a binary evaluation mechanism, assigning a value of 1 when the predicted answer matches the ground truth and 0 otherwise.
Format Reward in Text Generation.  To encourage ULMs to generate structured and interpretable textual responses, we adopt the format reward [18], which requires the model to enclose its thinking process inside <think> $\cdots$ </think>, and provide its final answer within <answer> and <answer> tags. The format reward $\mathcal{R}_{\text{Format}}$ returns 1 for strict compliance, and 0 otherwise.

### v3

@@ 1 Introduction
As large foundation models (LFMs) continue to advance in their general capabilities and breadth of knowledge, post-training [kumar2025llm, lv2025towards, tie2025survey, tang2025video, zhou2025reinforced] has emerged as a critical paradigm for further refining pretrained LFMs toward specialized applications, thereby facilitating task adaptation and human-aligned behaviors. Recently, reinforcement learning (RL)-based approaches [jaech2024openai, jaech2025openaio3, team2025kimi, yu2024rlhf, tan2025reason, rafailov2023direct, shao2024deepseekmath] have exhibited considerable promise due to their data efficiency and strong alignment abilities. A notable exemplar is DeepSeek-R1 [guo2025deepseek], which demonstrates that RL with verifiable rewards and the group relative policy optimization (GRPO) algorithm constitutes a practical and stable strategy that sidesteps explicit preference modeling [wang2024mdpo] and reward model learning [wang2025unified]. This promising paradigm indicates the significant potential of LFMs to acquire advanced capabilities and generalize effectively without dependence on large-scale, high-quality supervised data.
@@
In the multimodal AI research community, the prevailing implementation [liu2025seg, huang2025vision, deng2025boosting, peng2025lmm, liu2025visual, zhang2025r1vl, zhan2025vision, tan2025reason] of the GRPO algorithm centers on crafting diverse rule-based reward mechanisms to incentivize long-chain reasoning capabilities of multimodal large language models (MLLMs). These initiatives primarily target multimodal understanding, with a particular focus on visual and mathematical reasoning tasks. Conversely, its application to visual generation remains surprisingly limited, with only pioneering explorations [wang2025simplear, jiang2025t2i] suggesting its feasibility. More importantly, extending GRPO to unified MLLMs (ULMs) [chen2025januspro, xie2025show, wu2025vila, liu2025world] capable of concurrently performing visual understanding and generation tasks remains considerably under-explored. Intuitively, ULMs could significantly benefit from GRPO owing to their inherent advantages of cross-task synergy and LLM sharing, which enables ULMs to share reward signals across various tasks and effectively mitigate reward imbalance, particularly as GRPO operates by jointly ranking outputs within task-agnostic groups.
@@
This paper aims to enhance the understanding and generation capabilities of ULMs without relying on supervised data. We begin with a set of pilot experiments to explore efficient reinforcement learning paradigms. Specifically, we systematically examine four rule-based training strategies: (i) separate RL for individual tasks, (ii) separate RL with weight merging, (iii) cycle RL alternating between tasks, and (iv) unified RL with joint optimization. Our explorations reveal two critical findings. First, direct task-specific RL fails to achieve the anticipated improvements, particularly in visual generation, and even impairs other abilities. Second, compared with alternative strategies, unified RL showcases comprehensive advantages across tasks. These results demonstrate the synergistic co-evolution of dual capabilities under a shared policy optimization paradigm.
@@
In light of our preliminary findings, we propose CoRL, a co-reinforcement learning framework designed to synergistically improve the understanding and generation capabilities of ULMs. Specifically, CoRL follows a Foundation-then-Specialization paradigm and is implemented through a two-stage RL procedure: a unified RL stage for joint optimization of dual capabilities and a refined RL stage for task-specific enhancement. In the first stage, the policy ULM is optimized through a unified GRPO algorithm with diverse rewards on a carefully curated dataset spanning both understanding and generation tasks. To effectively guide policy optimization in visual generation, we introduce a bidirectional cycle consistency reward and a text-image matching reward, which together promote semantic consistency and faithfulness between synthesized images and their corresponding prompts. The designed rewards complement typical multimodal understanding rewards (i.e., accuracy and format) within a unified group, enabling cross-task joint optimization. In the subsequent stage, we independently reinforce the policy’s understanding and generation capabilities using respective rewards and tailored datasets for task-specific refinement.
@@
Applying the two-stage CoRL procedure to the baseline ULM Janus-Pro [chen2025januspro] yields ULM-R1, a unified model with reinforced capabilities in both understanding and generation. To comprehensively assess its performance, we conduct extensive comparisons against state-of-the-art unified MLLMs and dedicated models across both three visual generation and nine multimodal understanding benchmarks. Notably, ULM-R1 achieves substantial gains over its baseline on complex mathematical and logical reasoning tasks, such as WeMath (+15.2) and LogicVista (+10.6). These results underscore the effectiveness of CoRL, providing compelling empirical evidence for the efficacy of RL in jointly advancing visual understanding and generation tasks.
@@ 2 Related Work
Unified Multimodal Understanding and Generation.  Recent advancements [wu2025harmonizing, chen2025januspro, liu2025world, wu2025vila, wu2024liquid, team2024chameleon, xie2025show, zhou2024transfusion, xie2024muse, wang2024emu3, sun2024autoregressive, wang2024illume] have witnessed increasing attention to jointly model multimodal understanding and visual generation within a unified model. Pioneering attempts [ge2024seed, dong2024dreamllm] predominantly rely on continuous diffusion models, integrating external diffusion decoders for image synthesis. Inspired by autoregressive next-token prediction, a growing line of research [team2024chameleon, wu2025harmonizing, wang2024emu3, wu2024janus, chen2025januspro, qu2024tokenflow, liu2025world, wu2025vila, wu2024liquid, li2024synergen, ma2025unitok, zou2025omnimamba] encode visual inputs into discrete tokens and generate images in a fully autoregressive (F-AR) manner. Specifically, this approach employs a vector quantized (VQ) tokenizer [esser2021taming, yu2021vector] to convert images into discrete tokens, analogous to text tokenization. To mitigate information loss in VQ discretization, another stream of work [xie2025show, wang2024illume, huang2025illume+, kou2024orthus, fan2025unified, ma2024janusflow, yang2025hermesflow, chen2025blip3, tong2024metamorph, shi2024lmfusion] explores autoregressive and diffusion (AR-Diff) hybrid modeling approaches. Architecturally, these models typically comprise a vision autoencoder, a text tokenizer, and an LLM. Given the unified advantage of the F-AR model in generation manner, this work builds upon it to develop our co-reinforcement learning framework.
@@
RL-based Post-Training for MLLMs.  Post-training [zhou2025reinforced] aims to further enhance the performance of pretrained models for customized applications and user needs. Recently, RL [wiering2012reinforcement, wang2016learning] has emerged as a powerful post-training technique, enabling models to learn from feedback and align with human values. RL in MLLMs can be broadly categorized into two paradigms: (1) RL from human/AI feedback (RLHF) [wang2024emu3, yang2025hermesflow, wang2025unified, sun2023aligning, yu2024rlhf, yu2024rlaif, zhou2024aligning, li2024vlfeedback, zhao2023beyond, pi2024strengthening, wang2024mdpo, sarkar2024mitigating, wang2024enhancing, zhai2024fine] and (2) RL with verifiable reward mechanisms [wang2025simplear, zhan2025vision, tan2025reason, liu2025visual, liu2025seg, li2025videochat]. RLHF involves learning reward models from preference data before RL optimization, whereas the latter directly optimizes models using task-specific reward functions, bypassing explicit preference modeling. For example, DPO [rafailov2023direct] is a notable implementation of RLHF and has been adopted by Emu3 [wang2024emu3] and HermesFlow [yang2025hermesflow] to narrow the performance gap between understanding and generation. In contrast, GRPO [shao2024deepseekmath] exemplifies the second paradigm, simplifying reward formulation via group-wise relative advantage estimation. Our work also falls into this paradigm but diverges from prior work such as SimpleAR [wang2025simplear], which utilizes GRPO with external CLIP reward for autoregressive visual generation, and R1-like MLLMs [zhan2025vision, huang2025vision, tan2025reason, liu2025visual] that focus on incentivizing reasoning capabilities. First, our work demonstrates the significant potential of RL in co-optimizing understanding and generation, thereby broadening its applicability beyond reasoning. Moreover, we identify semantic consistency rewards and a co-evolutionary reinforcement strategy as crucial components in enhancing ULMs.
@@ 3.1 Preliminary
Group relative policy optimization (GRPO) [shao2024deepseekmath] is a value-free policy optimization algorithm with improved training stability and sample efficiency. Building upon PPO [schulman2017proximal], GRPO introduces a group-wise relative advantage approach to bound policy updates while maintaining optimization flexibility. Let $\pi_{\bm{\theta}}$ denote a policy parameterized by $\bm{\theta}$. Formally, given an input content $c$, the algorithm first samples a group of $G$ outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the current policy $\pi_{\bm{\theta}_{\mathrm{old}}}$. Each output is then evaluated using predefined, verifiable reward functions, yielding the reward set $\{r_{1},r_{2},\dots,r_{G}\}$. These rewards are subsequently normalized to compute group-relative advantages as follows:
@@ 3.2 Pilot Exploration
Figure 1: Results of different RL paradigms. Janus-Pro-1B [chen2025januspro] serves as the baseline.
Given the exceptional performance and data efficiency of DeepSeek-R1-Zero [guo2025deepseek], we explore the potential of ULMs to enhance understanding and generation capabilities without dependence on task-specific supervised fine-tuning. To accomplish this, we curate a dataset [^1] https://huggingface.co/datasets/mm-vl/x2x_rft_16k comprising 16K samples sourced from the COCO 2017 training split [lin2014microsoft]. Each sample includes a real image, an associated caption as a textual prompt for visual generation, and a corresponding QA pair for the multimodal understanding task. We adopt CLIP Score [radford2021learning] as the verifiable reward for image generation, along with a combination of formatting correctness and answer accuracy as the reward for text generation. We investigate four distinct RL paradigms: (i) separate RL, where understanding and generation tasks are independently optimized with their respective reward mechanisms; (ii) separate RL followed by weight merging, where each task is separately optimized, and the resulting weights are subsequently merged using a Gaussian distribution-based merging strategy [si2025unveiling] to incorporate both abilities; (iii) cycle RL, which employs a scheduled alternation between the two tasks throughout the training process; and (iv) unified RL, in which both tasks are jointly optimized within a unified paradigm to promote the co-evolution of dual capabilities.
As presented in Figure˜1, we observe that (1) direct task-specific RL fails to achieve the expected improvements for ULMs, particularly in the visual generation task, and may even impair performance on the other task; and (2) unified RL demonstrates substantial advantages over alternative paradigms. These findings indicate that the dual capabilities that co-evolve within a shared training framework contribute to enhanced cross-task synergy and knowledge transfer.
@@ 3.3.1 Verifiable Reward for Multimodal modeling
Bidirectional Cycle Consistency Reward in Text-to-Image Generation.  To encourage ULMs to generate images that faithfully depict the concepts and entities described in the input prompt, we introduce a bidirectional cycle consistency reward $\mathcal{R}_{\text{cycle}}$, which measures the consistency between predictions and ground truth in both visual and textual spaces. For visual consistency, we adopt LPIPS [zhang2018perceptual] to assess the patch-level perceptual similarity between the real image $\mathcal{I}_{\text{real}}$ and the synthesized image $\mathcal{I}_{\text{gen}}$. Textual consistency is implemented in a re-captioning manner. Specifically, we first employ BLIP [li2023blip] to generate a caption $\mathcal{C}_{\text{re-cap}}$ for each synthesized image, and then compute the SPICE [anderson2016spice] score between $\mathcal{C}_{\text{re-cap}}$ and its original prompt $\mathcal{P}_{\text{org}}$ to measure semantic fidelity. The combined bidirectional cycle reward is defined as:
@@
This bidirectional reward forms a closed feedback loop that promotes mutual consistency between texts and images, effectively penalizing hallucinated content and reinforcing prompt-aligned visual generation by simultaneously optimizing for both visual and textual consistency. Furthermore, $\mathcal{R}_{\text{cycle}}$ is normalized to the range [0, 1] before being combined to ensure that all rewards operate on comparable scales and to prevent any single component from dominating due to scale differences.
Text-Image Matching Reward.  While CLIP Score [radford2021learning] provides a holistic measure of text-image alignment, as shown in Sec.˜3.2, it underperforms due to its limited capacity for assessing fine-grained semantics. To address this limitation, we instead propose a text-image matching reward $\mathcal{R}_{\text{TIM}}$, which leverages the ULM itself to evaluate cross-modal alignment at the token level. Given a textual representation $\bm{T}=\{\bm{t}_{1},\bm{t}_{2},\ldots,\bm{t}_{L_{t}}\}\in\mathbb{R}^{L_{t}\times d}$ of the prompt and the corresponding visual representation $\bm{I}=\{\bm{i}_{1},\bm{i}_{2},\ldots,\bm{i}_{L_{i}}\}\in\mathbb{R}^{L_{i}\times d}$ of a generated image, the reward is computed as:
@@
where $L_{t}$ and $L_{i}$ are the sequence lengths of the textual and visual tokens, $d$ is the embedding dimension, and $\mathcal{R}_{\text{TIM}}$ is also be normalized to the range [0, 1]. This reward captures the fine-grained correspondence between textual concepts and visual elements through maximum cosine similarity, ensuring mutual alignment between visual tokens and their most relevant textual counterparts.
Accuracy Reward in Multimodal Question Answering.  Accuracy rewards leverage task-specific metrics to directly evaluate the correctness of ULM predictions. We consider two accuracy rewards tailored to different question types: $\mathcal{R}_{\text{MCQ-Acc}}$ for multi-choice questions and $\mathcal{R}_{\text{OE-Acc}}$ for open-ended questions. These rewards follow a binary evaluation mechanism, assigning a value of 1 when the predicted answer (i.e., the final answer parsed from within <answer> and </answer> tags) matches the ground truth and 0 otherwise.
Format Reward in Text Generation.  To encourage ULMs to generate structured and interpretable textual responses, we adopt the format reward [guo2025deepseek], which requires the model to enclose its thinking process inside <think> $\cdots$ </think>, and provide its final answer within <answer> and </answer> tags. The format reward $\mathcal{R}_{\text{Format}}$ returns 1 for strict compliance and 0 otherwise.